# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v3

@@ VII Conclusion
We have introduced a novel, camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera that is placed on the AMR at low height, and a laser crosshair that is projected from the AMR onto the chessboard floor and observed by the camera for position and heading estimation. The prototype achieved low mm localization and sub-degree heading estimation accuracy.
Our prototype will be used in future work to benchmark our radar-based indoor localization system. We have further discussed the many advantages this localization approach brings and how it may be improved and extended in various directions in future work.

### v4

@@ VII Conclusion
We have introduced a novel, camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera placed at low height on the AMR, and a laser diode that projects a crosshair from the AMR onto the chessboard floor. The crosshair on the floor is continuously observed by the camera for position and heading estimation. Our prototype achieved low mm localization and sub-degree heading accuracy. We have further discussed the many advantages this localization approach brings and how it may be improved and extended in various directions in future work.